Find the slope between the points.
(8, -1) and (10, -7)

Answer by greenestamps(13215)



slope = rise/run

run = distance left to right
rise = distance up (positive) or down (negative)

From (8,-1) to (10,-7) the run is 2 (from 8 to 10)

From (8,-1) to (10,-7) the rise is -6 (from -1 to -7)

The slope is rise/run = -6/2 = -3